OpenStudy (anonymous):

Solve |2x + 8| = 6. Show its equivalent compound sentence and show its solution please?

OpenStudy (anonymous):

-6<2x+8<6 =-6-8<2x+8-8<6-8 =-14<2x<-2 =-7<x<-1 So the value of x is either -7 or -1.

OpenStudy (anonymous):

thanks your really helping me out